OpenStudy (abmon98):

A cyclist is travelling along a straight road. The (t,v) graph above, where t is measured in seconds and v is measured in vms^-1, consists of three line segments and models the first 250m/s of the cylists journey. The cyclist has a constant speed of Vms^-1 in the interval 60

OpenStudy (abmon98):

end of the 250 seconds

OpenStudy (abmon98):



OpenStudy (abmon98):

i solved a and b im stucks in c

OpenStudy (abmon98):

stuck**

ganeshie8 (ganeshie8):

a) find an expression in terms of v for the distance travelled in meters in first 200 seconds

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

a) distance = area under velocity curve \(\large d = \frac{1}{2}60V + 100V + \frac{1}{2}40V\)

ganeshie8 (ganeshie8):

b) given that the average speed of the cylcist in the interval 0< or equal to t or < or equal to 200 is 6.75ms^-1 find V \(\large 200*6.75= \frac{1}{2}60V + 100V + \frac{1}{2}40V \) solve \(V\)

OpenStudy (abmon98):

V=9

ganeshie8 (ganeshie8):

yes i see ur work is looking good :)

OpenStudy (abmon98):

part c is too hard i cant figure out the v in time =250

ganeshie8 (ganeshie8):

from part b, we knw that he travelled 200*6.75 = 1350 meters in first 200 seconds

ganeshie8 (ganeshie8):

we just need to find out, how much distance he travelled between 200 and 250 seconds right ?

OpenStudy (abmon98):

yes